Two charges +q and +4q are separated a distance of 2r. Use Gauss’ law to find the electric
field at a distance r around the charge +question and 3r around the charge +q
1
Expert's answer
2019-07-24T15:43:19-0400

1) According to Gauss's law, the flux provided by the first and the second charge is


"\Phi_1=\frac{Q_1}{\epsilon_0}=\frac{q}{\epsilon_0},""\Phi_2=\frac{Q_2}{\epsilon_0}=\frac{4q}{\epsilon_0},"

and at a distance r between the charges the flux is


"\Phi=\Phi_1+\Phi_2=\frac{5q}{\epsilon_0}."

On the other hand, Gauss's law can be expressed in terms of electric field:

"E_1=\frac{\Phi_1}{4\pi r^2}=\frac{q}{4\pi\epsilon_0 r^2},""E_2=\frac{\Phi_2}{4\pi r^2}=\frac{4q}{4\pi\epsilon_0 r^2}."

Since the two charges are positive, in the middle between the charges the total field (directed from the larger charge) is


"E=E_2-E_1=\frac{3q}{4\pi\epsilon_0 r^2}."

2) Reasoning the same way, the flux around the charges will be as in the previous problem.


"\Phi_1=\frac{q}{\epsilon_0},""\Phi_2=\frac{4q}{\epsilon_0},"

and for distance of 3r from the smaller charge in direction of the largest charge the field will be


"E_1=\frac{\Phi_1}{4\pi\cdot9r^2}=\frac{q}{4\pi\epsilon_0\cdot9r^2},"

and since the charges are 2r from each other, that distance will be r behind the larger charge:


"E_2=\frac{\Phi_2}{4\pi\cdot r^2}=\frac{4q}{4\pi\epsilon_0\cdot r^2},"

the total electric field


"E=E_1+E_2=\frac{37q}{4\pi\epsilon_0\cdot 9r^2}."